OpenStudy (anonymous):

Help with logs?Cant use a calculator?:( Evaluate, without using a calculator 5^(-2)

OpenStudy (anonymous):

Oh wait i put the wrong one sowwy :( \[\log _{6}1\]

OpenStudy (anonymous):

log(a)y = a^y

OpenStudy (anonymous):

im a little confused.....would it be \[6^{1}\] than?

OpenStudy (anonymous):

yes :)

OpenStudy (anonymous):

ohh okay thankyou(: um may you help me with another one? Its a fraction?

OpenStudy (anonymous):

yesm

OpenStudy (anonymous):

thankyou!(: okay here it is\[\log _{6}\frac{ 1 }{ \sqrt[5]{36}}\] sorry it took a while to type it lol

OpenStudy (anonymous):

i did the first one wrong it should be log(a)x = y and a^y =x

OpenStudy (anonymous):

ohh... okay im confused now :(

OpenStudy (anonymous):

the second one is a little out of my league lol you should open a new question with that second part!

OpenStudy (anonymous):

lol alright but may you explain the first one pleaseee?(: <3

OpenStudy (anonymous):

so log(a)x is the exponent to which the base a must be raised to give x

OpenStudy (anonymous):

so to raise 6 by a power of zero will give an output of 1. any number raised to zero will give 1

OpenStudy (anonymous):

x^0=1

OpenStudy (anonymous):

2^0=1 and so on

OpenStudy (anonymous):

so.... okay im not 100% sure but would the answer be 1?

OpenStudy (anonymous):

Oh and another question (sorry lol) so do you always raise the little number by 0?

OpenStudy (anonymous):

log(6)1=0 6^0=1

OpenStudy (anonymous):

nope you 'exponentiate' the little number

OpenStudy (anonymous):

log(6)1=0 since 6^0=1

OpenStudy (anonymous):

okay okay...what does exponentiate mean im sorry for all these questions

OpenStudy (anonymous):

The act of raising a quantity to a power :)

OpenStudy (anonymous):

okay(: but im still confused :'(

OpenStudy (anonymous):

they confuse me too hehe

OpenStudy (anonymous):

log(a)x = y <=> a^y = x

OpenStudy (anonymous):

just remember that a is your base, the base will always be the same for any log

OpenStudy (anonymous):

any number raised to a power like a^x, a will be your base

OpenStudy (anonymous):

imm off gluck

OpenStudy (anonymous):

alright:/ thanks though <3